OpenStudy (anonymous):

Stats: Am I on the right path? The problem is, "Show that the variance of the finite population (c_1,...c_N) can be written as ...

OpenStudy (anonymous):

\[\sigma ^{2} = \frac{ { \sum_{i=1}^{N} }{ }c_i ^{2} }{ N } -\mu ^{2}\]

OpenStudy (anonymous):

So, my thought thus far has been, start with the formula for sigma^2\[\sigma ^{2} = \frac{ 1 }{ N } \sum_{i=1}^{N} (c_i-\mu)^{2}\] and then use the formula for mu instead of just the symbol for it.

hartnn (hartnn):

ok, those must be equally likely events to be able to use that formula ... ?

OpenStudy (anonymous):

its a weird varaition of it in my book, gimme 2 secs, ill take a picture.

OpenStudy (anonymous):

Its problem 8.15. I assume this is just using the formulae for mu and sigma^2 and shuffling things around until it looks like that, but I was wondering if I could get some confirmation of that before I spend an hour on a fruitless pursuit

hartnn (hartnn):

yeah, the end result is correct, and use \(\large \mu =\dfrac{1}{N}\sum \limits_{i=1}^Nx_i\)

OpenStudy (anonymous):

Ok perfect, Im probably about halfway through it, but I move frustratingly slow and wanted to make sure that was the right idea. Thanks!

hartnn (hartnn):

thank wiki though :P http://en.wikipedia.org/wiki/Variance

hartnn (hartnn):

i got the answer in 2 steps, \(\large= ...-2\mu \times \mu+\mu^2\ \\ \large = ...-\mu^2\) last steps....